OpenStudy (anonymous):

graph y=2x² 1 (parabola)

OpenStudy (anonymous):

complete the equation is that + or - 1 on the end

OpenStudy (anonymous):

oh my bad is +1

OpenStudy (anonymous):

ok graph opens upward

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

axis of symmetry is x = -b/2a= -0/4 = 0

OpenStudy (anonymous):

vertex is where x = 0 y = 1

OpenStudy (anonymous):